# Sparse is Enough in Fine-tuning Pre-trained Large Language Model

### v1

@@ 5.1 GLUE Benchmark
5.2 Instruction-tuning
(Wei et al., 2022) proposes to perform supervised fine-tuning on top of pre-trained language models to make language models align instructions and generate more meaningful content. (Taori et al., 2023) fine-tunes the pre-trained Llama on a 52k instruction-following dataset (Alpaca), which is generated by the technique introduced in (Wang et al., 2023). We fine-tune llama with Alpaca in different ways including full fine-tuning, LoRA and our SIFT. We evaluate the language model on multiple benchmarks including MMLU and HumanEval. For MMLU, the evaluation metric is the accuracy of the generating answer in zero-shot setting, while the code evaluation task HumanEval is evaluated by the pass@1 and pass@10 metric. Higher values are better for all metrics. Detailed experimental hyper-parameters are described in the Appendix. Table 3 demonstrates that sparse structures like SIFT can inject knowledge into the pre-trained large language models and exhibit competitive performance.
@@ 5.3 Further Analysis
In this subsection, we primarily conduct further experimental analysis towards SIFT. This involves comparing the performance differences when applying SIFT on different parameters, as well as the differences between SIFT’s gradient-based component selection and random selection. Finally, we discuss the impact of the sparsity rate on the performance of downstream tasks.
5.3.1 SIFT Vs Random
SIFT selects the component with the highest absolute gradient value in the first batch for the subsequent updates. A natural question arises: does the selection method of SIFT have an advantage over random selection? Additionally, in previous experiments, we heuristically choose the $W_{q},W_{k},W_{v}$ and $W_{o}$ in the attention modules for sparse updating. Under the condition of maintaining the same number of trainable parameters, including more modules to apply SIFT means that each module has fewer components available for updates. We want to know whether having fewer modules with more components or more modules with fewer components performs better in the fine-tuning process. In Table 4, we compare the performance of SIFT and random selection in updating different modules. We choose two small datasets from the GLUE benchmark, RTE and MRPC, for evaluation. The final performances do not show a consistent conclusion in the choice of different update modules. Therefore, balancing the number of components and modules, we choose to update $W_{q},W_{k},W_{v}$ and $W_{o}$ as our final choice. Compared to random selection, SIFT’s gradient-based selection method shows a clear advantage in the final performance of downstream tasks. However, unexpectedly, although not as effective as SIFT in terms of performances, sparse updates in randomly selected components can still effectively adapt pre-trained models to downstream tasks, with only a small performance gap. One hypothesis is that the property of sparsity is broadly present in pre-trained models, with quasi-sparse gradients being just one of its important representations, which still requires further research.
@@ 5.3.2 Sparsity Rate Analysis
We test the efficiency of different methods, that is, the trend of performance with the number of trainable parameters. We compare SIFT, Random selection and LoRA with varied numbers of trainable parameters on the MRPC task. The results show that SIFT is able to adapt the downstream task more efficiently with smaller parameters, while the performance gap between SIFT and LoRA is smoothed out as the number of parameters increases. Consistent with the conclusion of 5.3.1, we find that when the number of parameters reaches a certain threshold, Random is also able to adapt the downstream task, but it is significantly inferior to SIFT and LoRA in terms of the performances for small parameters.
@@ 6 Conclusion
In this paper, we reveal the gradient properties of pre-trained language models on downstream domains. By visualizing the loss landscape and the gradient distribution, we find that the gradients of pre-trained models exhibit a sparse-like property, which we refer to as the Quasi-Sparse gradient distribution. Based on this, we introduce a component-sparse fine-tuning paradigm named SIFT, which differs from previous parameter-efficient fine-tuning methods that involve injecting small plug-ins (such as Adapters and LoRA) or module-sparsity (such as Bias-only), and propose a memory-efficient implementation. This method is a more streamlined approach for fine-tuning, and experiments demonstrate that it is on par with the mainstream PEFT methods.
Limitation
Although we have discovered that the gradients of pre-trained models on downstream tasks exhibit a quasi-sparse nature, the specific reasons for this are not yet clear. How to utilize this feature to aid in better pre-training will be a viable direction. Meanwhile, the selection of modules to update remains heuristic, and whether the selection of modules and allocation of component quotas can also be done through gradients is a question worth researching. In this paper, we present a memory-efficient implementation of SIFT but It must be acknowledged that our implementation currently only supports training on a single card and that the training speed significantly decreases with the increase of the sparsity rate. How to implement SIFT more efficiently remains an unresolved issue.

### v3

@@ 5.3 Further Analysis
We primarily conduct further experimental analysis towards SIFT involving comparison of performance differences when applying SIFT on different parameters, as well as the differences between SIFT gradient-based component selection and random selection. Finally, we discuss the impact of the sparsity rate on the performance of downstream tasks.
5.3.1 SIFT Vs. Random
@@
SIFT selects the component with the highest absolute gradient value in the first batch for the subsequent updates. A natural question arises: does this selection have an advantage over random selection? Additionally, in previous experiments, we heuristically choose the $W_{q},W_{k},W_{v}$ and $W_{o}$ in the attention modules for sparse updating. Under the condition of maintaining the same number of trainable parameters, including more modules to apply SIFT means that each module has fewer trainable components. We want to know whether having fewer modules with more components or more modules with fewer components performs better in the fine-tuning process. In Table 4, we compare the performance of SIFT and random selection in updating different modules. We choose two small datasets from the GLUE benchmark, RTE and MRPC, for evaluation. The final performances do not show a consistent conclusion in the choice of different update modules. Therefore, balancing the number of components and modules, we choose to update $W_{q},W_{k},W_{v}$ and $W_{o}$ as our final choice. Compared to random selection, SIFT gradient-based selection method shows a clear advantage in the final performance of downstream tasks. However, although not as effective as SIFT in terms of performances, sparse updates in randomly selected components can still effectively adapt pre-trained models to downstream tasks, with only a small performance gap. This further confirms that due to the shift of the prior induced by pre-training, the model is away from non-expressive submanifolds in parameter space, resulting in redundancy across dimensions. Searching along partial dimensions is sufficient to find a relatively good solution.
@@ 5.3.2 Sparsity Rate Analysis
We investigate the efficiency of different methods, that is, the trend of performance with the number of trainable parameters (budget). We compare performances of SIFT, Random selection and LoRA with varied budgets on MRPC dataset. Figure 6 show that SIFT is able to adapt the downstream task more efficiently in a less budget, while the performance gap between SIFT and LoRA is smoothed out as the budget increases. Consistent with the conclusion in Section 5.3.1, we find that when the number of parameters reaches a certain threshold, Random is also able to adapt the downstream task, but it is significantly inferior to SIFT and LoRA in terms of performances in less budgets.
@@ 6 Conclusion
In this paper, we adopt PAC-Bayesian generalization error bounds to analyze the pre-training-fine-tuning paradigm and attribute the performance differences to the prior distribution shift. We verify the shift of the prior through investigating differences of pre-trained models in loss landscapes and gradient distributions. Pre-trained models, compared to models trained from scratch, exhibit sharper oscillations in the loss landscape and gradients display a quasi-sparse property, where partial components dominates the majority of the gradient norm. Due to these differences, we demonstrate that the search space dimensions required for pre-trained models to reach optimal solutions during fine-tuning are compressed, allowing effective adaptation to downstream tasks by fine-tuning only a small portion of parameters. Based on this, we propose a component-sparse fine-tuning strategy, named SIFT. We validate the effectiveness of using SIFT for fine-tuning on varied tasks including GLUE Benchmark and instruction-tuning for large language models. Furthermore, we have further discussions on SIFT, highlighting its efficiency from the comparison of different component selection methods and the performance trends of different methods as it varies with the number of trainable parameters. Overall, we demonstrates the feasibility of fine-tuning language models entirely in a component-sparse way.
Impact Statement
The main motivation of this work has been to design a simple but efficient fine-tuning method with a better understanding of pre-trained models. By uncovering certain important features and properties within pre-trained models, such as the quasi-sparsity of gradients proposed in this paper, we aim to utilize resources more efficiently during the fine-tuning process. This work can be of interest to researchers focusing on efficient fine-tuning and sparsity in deep models. There is no foreseeable negative societal/ethical impact of our work at this time.
Appendix A A Primer on PAC-Bayesian Generalization Error Bounds.
Due to limited pages of the main paper, all statements regarding PAC-Bayesian theory in the main text are kept as concise as possible. Therefore, in this section, we primarily address readers unfamiliar with PAC Learning or PAC-Bayesian Theorems. We will rigorously restate the theorems mentioned in the main paper, providing proofs for the simpler theorems and referencing related papers for the complex one.
The theorems in this section are introduced in previously published works. The presented overview here serves as a reminder and provides references for the readers as needed. The notation in this section remains consistent with that in the main paper.
Theorem A.1.
If $\mathcal{H}$ is a finite hypothesis space, for any hypotheses $h\in\mathcal{H}$, any loss functions $l$ bounded in $[0,1]$, $10<\delta<$ 1, with a probability at least $1-\delta$ over the selection of n i.i.d. samples,
^$\hat{\mathcal{R}}(h)\leq\mathcal{R}(h)+\sqrt{\frac{log(|\mathcal{H}|)+log(1/\delta)}{2n}}$
Proof.
Using Hoeffding’s inequality, for any $\epsilon>0$,
$\mathbb{P}\{\Delta(h)=\hat{\mathcal{R}}(h)-\mathcal{R}(h)\geq\epsilon\}\leq e^{-2n\epsilon^{2}}$
Applying the Union Bound,
$\mathbb{P}\{\exists h\in\mathcal{H},\Delta(h)>\epsilon\}\leq\underset{h\in\mathcal{H}}{\sum}\mathbb{P}\{\Delta(h)>\epsilon\}\leq|\mathcal{H}|e^{-2n\epsilon^{2}}$
Let $|\mathcal{H}|e^{-2n\epsilon^{2}}=\delta$, and solve the equation for $\epsilon$, we have $\epsilon=\sqrt{\frac{log(|\mathcal{H}|)+log(1/\delta)}{2n}}$.
Theorem A.2.
(McAllester, 1998). If $\mathcal{H}$ is a finite hypothesis space, for any probability distribution $\mathcal{P}$ over $\mathcal{H}$ that assigns non-zero values, any hypotheses $h\in\mathcal{H}$, any loss functions $l$ bounded in $[0,1]$, $10<\delta<$ 1, with a probability at least $1-\delta$ over the selection of n i.i.d. samples,
^$\hat{\mathcal{R}}(h)\leq\mathcal{R}(h)+\sqrt{\frac{log(1/\mathcal{P}(h))+log(1/\delta)}{2n}}$
Proof.
For any fixed $h\in\mathcal{H}$, considering the Chernoff Bound, we have
$\mathbb{P}\{\Delta(h)>\sqrt{\frac{log(1/\mathcal{P}(h))+log(1/\delta)}{2n}}\}=\mathbb{P}\{\Delta^{2}(h)>\frac{log(1/\mathcal{P}(h))+log(1/\delta)}{2n}\}\leq\mathcal{P}(h)\delta$
Then using the Union Bound,
$\mathbb{P}\{\exists h\in\mathcal{H},\Delta(h)>\sqrt{\frac{log(1/\mathcal{P}(h))+log(1/\delta)}{2n}}\}\leq\underset{h\in\mathcal{H}}{\sum}\mathbb{P}\{\Delta(h)>\sqrt{\frac{log(1/\mathcal{P}(h))+log(1/\delta)}{2n}}\}\leq\underset{h\in\mathcal{H}}{\sum}\mathcal{P}(h)\delta=\delta$
Theorem A.3.
(McAllester, 2003). For any probability distribution $\mathcal{P}$ on a possibly uncountable hypothesis space $\mathcal{H}$ and any measurable loss function $l$, with a probability at least $1-\delta$,
$\forall\mathcal{Q}\ on\ \mathcal{H},\underset{h\sim\mathcal{Q}}{\mathbb{E}}[R(h)]\leq\underset{h\sim\mathcal{Q}}{\mathbb{E}}[\hat{R}(h)]+\sqrt{\frac{\mathbb{KL}(Q||P)+\log({n/\delta})+2}{2n-1}}$
A weak version of Theorem A.3 is first introduced in (McAllester, 1999). The complete proof of Theorem A.3 is in (McAllester, 2003). (Maurer, 2004; Catoni, 2007; Thiemann et al., 2017) have developed refinements on it.
Appendix B Experimental Details
B.1 GLUE Benchmark Setting
The table 5 shows our hyper-parameters settings on the GLUE benchmark experiments.
Table 5: The hyperparameters of the GLUE Benchmark experiments
Method Dataset MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
Optimizer AdamW
Warmup Ratio 0.06
LR Schedule Linear
SIFT Batch size 32
Epochs 10 15 20 20 10 20 20 30
Learning rate 7e-5 7e-5 7e-5 7e-5 5e-5 7e-5 7e-5 8e-5
Weight decay 0.1
Max Seq. length 512
Sparsity rate 0.8%
SIFT modules $W_{q},W_{k},W_{v},W_{o}$
B.2 Instruction-tuning Setting
The table 6 shows our hyper-parameters settings on the Instruction-tuning experiments.
Table 6: The hyper-parameters of the Instruction-tuning experiments
Method Model Llama-7b Llama-13b Llama-33b
Optimizer AdamW
Warmup Ratio 0.03
LR Schedule Linear
Full Batch size 128
Epochs 3
Learning rate 2e-5
Weight decay 0.
Max Seq. length 2048
LoRA Batch size 128
Epochs 3
Learning rate 2e-4 2e-4 1e-4
Weight decay 0.
Max Seq. length 2048
LoRA rank $r_{q}=r_{v}=128$
LoRA alpha 256
SIFT Batch size 128
Epochs 3
Learning rate 2e-4 9e-5 8e-5
Weight decay 0.
Max Seq. length 2048
Sparsity rate 5% 4% 3 %
SIFT modules $W_{q},W_{k},W_{v},W_{o}$
B.3 Results of differernt categories in MMLU Benchmark
Figures 7, 8, and 9 serve as a supplement to Table 3, illustrating the score comparisons for various categories on MMLU benchmark under different configurations.
Figure 7: MMLU Evaluation results of Llama-7b fine-tuned by different methods
Figure 8: MMLU Evaluation results of Llama-13b fine-tuned by different methods
Figure 9: MMLU Evaluation results of Llama-33b fine-tuned by different methods
B.4 Generated Code Samples in HumanEval Benchmark
Table 7: Generated Code Samples.
Model Method Generated Samples Passed
Prompt
[…]
\"\"\" You’re given a list of deposit and withdrawal
[…]
zero balance. Your task is to detect if at any point
[…]
at that point function should return True.
[…]
"entry_point": "below_zero", "canonical_solution":
[…]
Llama-7B Vanilla
return operations[0] < 0
False
Full
[…]
return False
False
LoRA
[…]
False
SIFT
[…]
return False
True